# BaseCal: Unsupervised Confidence Calibration via Base Model Signals

### v1

@@ 1 Introduction
Hallucinations have become a critical challenge for large language models (LLMs). To mitigate this risk, a primary direction is to equip model responses with confidence scores that are well-calibrated with their actual accuracy (on_the_calibration; calibration_survey). Such confidence enables abstention from low-confidence answers to mitigate hallucinations or otherwise alert users to potential errors. However, widely adopted post-trained language models (PoLLMs) have been found to exhibit significant overconfidence, often assigning high confidence even to incorrect responses (achiam2023gpt; zhu-etal-2023-calibration).
@@
Considerable efforts have been devoted to mitigating this overconfidence. One direction involves supervised strategies, such as calibration-oriented fine-tuning (wang-etal-2025-towards-objective; xiao2025restoring) or temperature scaling using human annotation (on_the_calibration). However, these supervised methods rely heavily on human-labeled data, which is often difficult to obtain in real-world applications. In contrast, unsupervised approaches attempt to estimate confidence from PoLLM itself, such as aggregated token probabilities (malinin2021uncertainty), verbalized confidence (tian-etal-2023-just; xiong2024can), and sampling-based consistency (farquhar2024detecting; semantic_entropy). While alleviating the reliance on human labels, these methods remain constrained by the quality of PoLLM signals, which often retain a certain degree of miscalibration (tan-etal-2025-consistent; simhi-etal-2025-trust).
Inspired by findings that base LLMs are well-calibrated in multiple-choice tasks (luo2025your; xiao2025restoring), we investigate whether they can serve as an external reference to enhance PoLLM calibration in general QA tasks, i.e., free-form QA. Figure 1 (more results in §4.1) illustrates that base LLMs exhibit significantly superior confidence calibration compared to PoLLMs on the widely used QA dataset TriviaQA. Motivated by this observation, we seek to calibrate PoLLMs with their base counterparts.
@@
To mitigate this cost, we propose BaseCal-Proj, which approximates the base LLM’s confidence via a lightweight projection. Specifically, we extract the final-layer states of PoLLM and base LLM for the same input questions and PoLLM-generated responses, and train a projection module to map the state of PoLLM to the target base state. During inference, the PoLLM’s final-layer states are projected and passed through the base LLM’s output layer to approximate the base probability distribution. We then calculate the confidence by averaging the probabilities assigned to the PoLLM-generated tokens under this distribution. Since BaseCal-Proj involves only a lightweight projection and an output layer, it incurs negligible overhead compared to BaseCal-ReEval and sampling-based methods (semantic_entropy; farquhar2024detecting).
@@ 2.1 Overconfidence of LLMs
Emerging evidence indicates that post-training can induce systematic overconfidence, thus degrading model calibration (xiao2025restoring; leng2025taming; wang-etal-2025-towards-objective). This effect occurs in both instruction tuning and RLHF (zhu-etal-2023-calibration; leng2025taming), and strengthens with more tunable parameters (chen-etal-2023-close). Prior work attributes this phenomenon to factors such as data overlap between fine-tuning and pretraining corpora (wang-etal-2025-towards-objective), reward bias (leng2025taming), preference collapse (xiao2025restoring), and catastrophic forgetting (he2023preserving). To mitigate overconfidence, prior work has explored domain-specific fine-tuning (xiao2025restoring), confidence-aware reward (leng2025taming), and feature-preserving adaptation (he2023preserving).
@@ 2.2 Confidence Calibration for LLMs
Confidence calibration aims to align model confidence with the correctness of its predictions. On the one hand, supervised methods rely on human-labeled data to perform calibration, e.g., temperature scaling (on_the_calibration; xie2024calibrating; joy2023sample; yu2022robust) optimizes a single temperature parameter to rescale probabilities. Other methods fine-tune models to provide calibrated confidence (kapoor-etal-2024-calibration; tao-etal-2024-trust; chen-etal-2023-close). Despite their effectiveness, supervised methods suffer from both reliance on human-labeled data (shen2024thermometer) and limited generalization (liu2025on).
In contrast, unsupervised methods avoid reliance on human-annotated labels. Prompt-based methods directly query the model about the correctness of its own output, including P(true), which estimates confidence as the probability that the model judges its answer to be correct (kadavath2022languagemodelsmostlyknow), and verbalized confidence, where confidence is explicitly expressed in natural language (tian-etal-2023-just; zhang-etal-2024-calibrating; lin2022teaching; xiong2024can). Sampling-based methods estimate confidence from the uncertainty across response samples, either via semantic agreement (manakul-etal-2023-selfcheckgpt; xiong2024can; lin2024generating; raj2023semantic) or through semantic entropy (farquhar2024detecting; semantic_entropy; nikitin2024kernel). However, these methods remain limited by their reliance on signals from the PoLLMs themselves, which preserve some extent of overconfidence (tan-etal-2025-consistent; simhi-etal-2025-trust).
Closely related to our study, luo2025your optimizes a temperature parameter to rescale PoLLM’s output probabilities to match those of base LLMs. In contrast to their probability-level adjustments, we recover calibration from the model’s hidden states, thereby leveraging richer internal information. More importantly, while their method is restricted to multiple-choice formats, our approach is more general and natively supports free-form generation, which is the predominant interaction paradigm for current LLMs
@@ 3 Preliminaries
To quantify the degree of miscalibration, we employ the Expected Calibration Error (ECE) (on_the_calibration), which partitions the samples into $M$ equal-width confidence bins $\{B_{1},\dots,B_{M}\}$, then calculates the absolute gap between average accuracy $\text{acc}(B_{m})$ and average confidence $\text{conf}(B_{m})$:
@@
We primarily focus on free-form QA as it aligns better with the primary usage paradigm of current LLMs. In this task, the aggregated token probability (malinin2021uncertainty) is commonly treated as the Vanilla confidence, as it represents the most direct utilization of the model’s signal, without requiring specific prompt design (tian-etal-2023-just) or multiple sampling (farquhar2024detecting). We employ average aggregation following orgad2024llmsknowshowintrinsic to mitigate the effect of sequence length.
@@ 4.1 Motivation
Current unsupervised calibration methods are limited by the signals from PoLLMs, e.g., verbalized confidences are often very high (leng2025taming), and consistency can fail under self-consistent errors (tan-etal-2025-consistent). This motivates us to identify a more reliable external reference beyond PoLLMs. Inspired by findings that base LLMs are well-calibrated in multiple-choice tasks (luo2025your; xiao2025restoring; achiam2023gpt), we investigate whether this superior calibration persists in more challenging free-form QA tasks.
The experiments are conducted on the widely used QA dataset TriviaQA, using Qwen2.5 (yang2024qwen2), Llama3.1 (grattafiori2024llama3), and the fully open-source Olmo2 (olmo20242), which provide checkpoints at different post-training stages. This enables us to analyze the effects of various post-training strategies. As shown in Figure 2, PoLLMs are consistently more miscalibrated than their base counterparts across all three model families, indicating that the better calibration of base LLMs persists in free-form QA. Moreover, results on Olmo2 indicate that diverse post-training methods all impair confidence calibration, revealing a shared limitation of existing post-training approaches. Leveraging these insights, our core idea is to calibrate PoLLMs with their base counterparts, thereby eliminating the reliance on miscalibrated signals from the PoLLMs.
@@ 4.3 BaseCal-Proj
Table 1: Calibration results across four datasets and three PoLLMs. Bold indicates the best unsupervised calibration performance, while underlining denotes the second best. Since DACA is specifically designed for multiple-choice tasks, it is marked with “-” on free-form QA datasets.
@@ 5.1 Experimental Setup
We conduct experiments across a diverse set of model families, including Qwen2.5 (yang2024qwen2), Llama3.1 (grattafiori2024llama3), and Olmo2 (olmo20242) series. For each post-trained model, we use its corresponding pre-trained counterpart to perform calibration.
@@
Our evaluation covers an extensive range of free-form question answering benchmarks, including TriviaQA (triviaqa), Natural Questions (NQ; lee-etal-2019-latent), SQuAD (squad), and WebQuestions (WebQ; webquestion). The correctness of responses is evaluated using LLM-as-a-judge, following tian-etal-2023-just; orgad2024llmsknowshowintrinsic. We further validate the effectiveness of LLM-as-a-judge via human verification. Appendix A.2 shows more details. To compare with DACA, we also include the widely-adopted multiple-choice benchmark MMLU (mmlu), which spans 57 subjects from STEM to humanities. Appendix A.1 shows detailed statistics of these datasets.
@@
Following previous works (tian-etal-2023-just; kadavath2022languagemodelsmostlyknow), we evaluate calibration using ECE (on_the_calibration) as introduced in Section˜3 and the Brier Score (BS; brier), which measures the mean squared error between confidence and correctness labels.
@@
    Vanilla. This approach directly employs the model’s native probabilities as confidence. For free-form QA, we follow prior work (orgad2024llmsknowshowintrinsic; factualconfidence; malinin2021uncertainty) and aggregate the token-level probabilities of the generated sequences to obtain a single confidence score.
@@
    P(True). This method prompts the model to self-assess its own output and takes the predicted probability of answering “True” as the confidence score (kadavath2022languagemodelsmostlyknow).
@@
    Verbalization. This method prompts PoLLMs to express confidence in natural language. We use the prompt from tian-etal-2023-just.
@@
    Semantic Entropy (SE). SE (farquhar2024detecting; semantic_entropy) samples multiple responses and computes the entropy over semantic clusters derived from these samples. Following li-etal-2025-towards, we normalize the SE values to the range $[0,1]$.
@@
    DACA. DACA (luo2025your), which is designed specifically for the multiple-choice format, uses examples where the PoLLM and base LLM produce the same choice, and then learns a temperature to align the PoLLM’s probabilities with those of the pre-trained model.
Beyond unsupervised baselines, we include Temperature Scaling (TS; on_the_calibration) as a supervised reference baseline, which learns a temperature parameter by minimizing the negative log-likelihood with respect to answer correctness. More details are shown in Appendix A.3.
@@ 5.5 Impact of Different PoLLMs (RQ4)
Subsequently, the influence of model size is analyzed using Qwen2.5-7b/14b/32b/72b-instruct. Table 4 shows that BaseCal-Proj consistently yields substantial improvements across all scales. Furthermore, BaseCal-Proj achieves superior performance on larger models, due to the stronger calibration capabilities of larger base LLMs, a trend also observed in previous work (zhu-etal-2023-calibration).
@@ 5.6 Application (RQ5)
Selective classification (selective_classification) enables a model to abstain from making predictions when confidence is low, thereby enhancing reliability by trading off coverage for higher accuracy. This mechanism is particularly critical for LLMs in decision-making tasks, where unreliable outputs can lead to severe consequences. In this section, we evaluate the effectiveness of our proposed BaseCal-Proj in the context of selective classification. Figure˜6 compares the accuracy of the vanilla baseline and BaseCal-Proj by varying the confidence threshold from 0.5 to 0.95, where prediction below the threshold is rejected. Experimental results demonstrate that BaseCal-Proj consistently surpasses the vanilla method in accuracy across all confidence thresholds, highlighting its superior capability in identifying reliable predictions.
@@ A.2 Details about LLM-as-a-Judge
To assess the correctness of model outputs in QA tasks, we employ an LLM-as-a-judge to determine whether the generated responses are semantically equivalent to the ground truth, following tian-etal-2023-just; simpleqa. For reproducibility, we utilize the powerful open-source Qwen2.5-14B-Instruct model as our judge. Inspired by the evaluation protocol from simpleqa, we use the specific prompt detailed in Section˜A.5 to verify response correctness. To validate the quality of the LLM judgment, we conducted a manual audit on 100 randomly sampled instances. Comparative verification against human ground truth reveals a disagreement rate of only 1% with human annotations, confirming the reliability of our automated evaluation process.
@@ A.3 Further Information about Baselines
(1) Vanilla. Several studies have employed the aggregated token probabilities as the vanilla confidence (orgad2024llmsknowshowintrinsic; factualconfidence; malinin2021uncertainty). Following implementation in orgad2024llmsknowshowintrinsic, we average the probabilities of all tokens in a response as the confidence.
(2) Verbalization. This method leverages the instruction-following capability of models to express confidence in natural language. Following the implementation of tian-etal-2023-just, we employ the following prompt:
@@
(3) P(True). This method follows the prompting strategy introduced by kadavath2022languagemodelsmostlyknow, where the LLM is directly queried to assess the correctness of its own output. P(True) also requires an additional forward pass on the PoLLMs after the initial response is generated. Specifically, we construct the following prompt:
@@
(4) Semantic Entropy. As proposed by semantic_entropy; farquhar2024detecting, semantic entropy estimates confidence by sampling multiple responses for the same question and computing their semantic consistency. Following the implementation details recommended by semantic_entropy, we set the sampling parameters as follows: temperature 0.5, number of samples 10, top_p = 1.0, and top_k = -1. Since this metric was originally designed for hallucination detection, its raw scores do not fall within the $[0,1]$ interval. Therefore, following li-etal-2025-towards, we normalize the semantic entropy values to the range $[0,1]$ to facilitate the calculation of ECE.
@@
(5) DACA. DACA (luo2025your) is an unsupervised approach that optimizes a temperature parameter to scale the probability of PoLLM, thereby approximating the probability distribution of the base LLM. This probability-level scaling necessitates that both the base LLM and PoLLM produce the exact same top-1 token for a given input, a constraint theoretically proved and empirically analyzed in Section 3 (pp. 4-5) of luo2025your. This limitation restricts DACA’s applicability primarily to multiple-choice tasks, as identical token generation is very rare for base LLM and PoLLM in free-form QA. Accordingly, we compare our method with DACA on the MMLU benchmark in the main experiments. We utilize the official implementation (luo2025your) and optimize DACA using the same training set employed for our proposed method.
(6) Temperature Scaling. We directly adopt the implementation of temperature scaling from xia-etal-2025-influences. This is a supervised method that optimizes an optimal temperature parameter to rescale probabilities, thereby minimizing the Expected Calibration Error (ECE). The training process requires questions and human-labeled answers to determine response correctness. In our implementation, we utilize the entire training dataset to optimize the temperature.

### v2

@@ 1 Introduction
Hallucinations have become a critical challenge for large language models (LLMs). To mitigate this risk, a primary direction is to equip model responses with confidence scores that are well-calibrated with their actual accuracy (Guo et al., 2017; Geng et al., 2024). Such confidence enables abstention from low-confidence answers to mitigate hallucinations or otherwise alert users to potential errors. However, widely adopted post-trained language models (PoLLMs) have been found to exhibit significant overconfidence, often assigning high confidence even to incorrect responses (Achiam et al., 2023; Zhu et al., 2023).
@@
Considerable efforts have been devoted to mitigating this overconfidence. One direction involves supervised strategies, such as calibration-oriented fine-tuning (Wang et al., 2025; Xiao et al., 2025) or temperature scaling using human annotation (Guo et al., 2017). However, these supervised methods rely heavily on human-labeled data, which is often difficult to obtain in real-world applications. In contrast, unsupervised approaches attempt to estimate confidence from PoLLM itself, such as aggregated token probabilities (Malinin and Gales, 2021), verbalized confidence (Tian et al., 2023; Xiong et al., 2024), and sampling-based consistency (Farquhar et al., 2024; Kuhn et al., 2023). While alleviating the reliance on human labels, these methods remain constrained by the quality of PoLLM signals, which often retain a certain degree of miscalibration (Tan et al., 2025; Simhi et al., 2025).
Inspired by findings that base LLMs are well-calibrated in multiple-choice tasks (Luo et al., 2025; Xiao et al., 2025), we investigate whether they can serve as an external reference to enhance PoLLM calibration in general QA tasks, i.e., free-form QA. Figure 1 (more results in §4.1) illustrates that base LLMs exhibit significantly superior confidence calibration compared to PoLLMs on the widely used QA dataset TriviaQA. Motivated by this observation, we seek to calibrate PoLLMs with their base counterparts.
@@
To mitigate this cost, we propose BaseCal-Proj, which approximates the base LLM’s confidence via a lightweight projection. Specifically, we extract the final-layer states of PoLLM and base LLM for the same input questions and PoLLM-generated responses, and train a projection module to map the state of PoLLM to the target base state. During inference, the PoLLM’s final-layer states are projected and passed through the base LLM’s output layer to approximate the base probability distribution. We then calculate the confidence by averaging the probabilities assigned to the PoLLM-generated tokens under this distribution. Since BaseCal-Proj involves only a lightweight projection and an output layer, it incurs negligible overhead compared to BaseCal-ReEval and sampling-based methods (Kuhn et al., 2023; Farquhar et al., 2024).
@@ 2.1 Overconfidence of LLMs
Emerging evidence indicates that post-training can induce systematic overconfidence, thus degrading model calibration (Xiao et al., 2025; Leng et al., 2025; Wang et al., 2025). This effect occurs in both instruction tuning and RLHF (Zhu et al., 2023; Leng et al., 2025), and strengthens with more tunable parameters (Chen et al., 2023). Prior work attributes this phenomenon to factors such as data overlap between fine-tuning and pretraining corpora (Wang et al., 2025), reward bias (Leng et al., 2025), preference collapse (Xiao et al., 2025), and catastrophic forgetting (He et al., 2023). To mitigate overconfidence, prior work has explored domain-specific fine-tuning (Xiao et al., 2025), confidence-aware reward (Leng et al., 2025), and feature-preserving adaptation (He et al., 2023).
@@ 2.2 Confidence Calibration for LLMs
Confidence calibration aims to align model confidence with the correctness of its predictions. On the one hand, supervised methods rely on human-labeled data to perform calibration, e.g., temperature scaling (Guo et al., 2017; Xie et al., 2024; Joy et al., 2023; Yu et al., 2022) optimizes a single temperature parameter to rescale probabilities. Other methods fine-tune models to provide calibrated confidence (Kapoor et al., 2024; Tao et al., 2024; Chen et al., 2023). Despite their effectiveness, supervised methods suffer from both reliance on human-labeled data (Shen et al., 2024) and limited generalization (Liu et al., 2025).
In contrast, unsupervised methods avoid reliance on human-annotated labels. Prompt-based methods directly query the model about the correctness of its own output, including P(true), which estimates confidence as the probability that the model judges its answer to be correct (Kadavath et al., 2022), and verbalized confidence, where confidence is explicitly expressed in natural language (Tian et al., 2023; Zhang et al., 2024; Lin et al., 2022; Xiong et al., 2024). Sampling-based methods estimate confidence from the uncertainty across response samples, either via semantic agreement (Manakul et al., 2023; Xiong et al., 2024; Lin et al., 2024; Raj et al., 2025) or through semantic entropy (Farquhar et al., 2024; Kuhn et al., 2023; Nikitin et al., 2024). However, these methods remain limited by their reliance on signals from the PoLLMs themselves, which preserve some extent of overconfidence (Tan et al., 2025; Simhi et al., 2025).
Closely related to our study, Luo et al. (2025) optimizes a temperature parameter to rescale PoLLM’s output probabilities to match those of base LLMs. In contrast to their probability-level adjustments, we recover calibration from the model’s hidden states, thereby leveraging richer internal information. More importantly, while their method is restricted to multiple-choice formats, our approach is more general and natively supports free-form generation, which is the predominant interaction paradigm for current LLMs
@@ 3 Preliminaries
To quantify the degree of miscalibration, we employ the Expected Calibration Error (ECE) (Guo et al., 2017), which partitions the samples into $M$ equal-width confidence bins $\{B_{1},\dots,B_{M}\}$, then calculates the absolute gap between average accuracy $\text{acc}(B_{m})$ and average confidence $\text{conf}(B_{m})$:
@@
We primarily focus on free-form QA as it aligns better with the primary usage paradigm of current LLMs. In this task, the aggregated token probability (Malinin and Gales, 2021) is commonly treated as the Vanilla confidence, as it represents the most direct utilization of the model’s signal, without requiring specific prompt design (Tian et al., 2023) or multiple sampling (Farquhar et al., 2024). We employ average aggregation following Orgad et al. (2025) to mitigate the effect of sequence length.
@@ 4.1 Motivation
Current unsupervised calibration methods are limited by the signals from PoLLMs, e.g., verbalized confidences are often very high (Leng et al., 2025), and consistency can fail under self-consistent errors (Tan et al., 2025). This motivates us to identify a more reliable external reference beyond PoLLMs. Inspired by findings that base LLMs are well-calibrated in multiple-choice tasks (Luo et al., 2025; Xiao et al., 2025; Achiam et al., 2023), we investigate whether this superior calibration persists in more challenging free-form QA tasks.
The experiments are conducted on the widely used QA dataset TriviaQA, using Qwen2.5 (Yang et al., 2024), Llama3.1 (Grattafiori et al., 2024), and the fully open-source Olmo2 (Walsh et al., 2025), which provide checkpoints at different post-training stages. This enables us to analyze the effects of various post-training strategies. As shown in Figure 2, PoLLMs are consistently more miscalibrated than their base counterparts across all three model families, indicating that the better calibration of base LLMs persists in free-form QA. Moreover, results on Olmo2 indicate that diverse post-training methods all impair confidence calibration, revealing a shared limitation of existing post-training approaches. Leveraging these insights, our core idea is to calibrate PoLLMs with their base counterparts, thereby eliminating the reliance on miscalibrated signals from the PoLLMs.
@@ 4.3 BaseCal-Proj
Table 1: Calibration results across 5 datasets and 3 PoLLMs. Bold indicates the best unsupervised calibration performance, while underlining denotes the second best. Since DACA is specifically designed for multiple-choice tasks, it is marked with “-” on free-form QA datasets.
@@ 5.1 Experimental Setup
We conduct experiments across a diverse set of model families, including Qwen2.5 (Yang et al., 2024), Llama3.1 (Grattafiori et al., 2024), and Olmo2 (Walsh et al., 2025) series. For each post-trained model, we use its corresponding pre-trained counterpart to perform calibration.
@@
Our evaluation covers an extensive range of free-form question answering benchmarks, including TriviaQA (Joshi et al., 2017), Natural Questions (NQ; Lee et al., 2019), SQuAD (Rajpurkar et al., 2016), and WebQuestions (WebQ; Berant et al., 2013). The correctness of responses is evaluated using LLM-as-a-judge, following Tian et al. (2023); Orgad et al. (2025). We further validate the effectiveness of LLM-as-a-judge via human verification. Appendix A.2 shows more details. To compare with DACA, we also include the widely-adopted multiple-choice benchmark MMLU (Hendrycks et al., 2021), which spans 57 subjects from STEM to humanities. Appendix A.1 shows detailed statistics of these datasets.
@@
Following previous works (Tian et al., 2023; Kadavath et al., 2022), we evaluate calibration using ECE (Guo et al., 2017) as introduced in Section˜3 and the Brier Score (BS; BRIER, 1950), which measures the mean squared error between confidence and correctness labels.
@@
    Vanilla. This approach directly employs the model’s native probabilities as confidence. For free-form QA, we follow prior work (Orgad et al., 2025; Mahaut et al., 2024; Malinin and Gales, 2021) and aggregate the token-level probabilities of the generated sequences to obtain a single confidence score.
@@
    P(True). This method prompts the model to self-assess its own output and takes the predicted probability of answering “True” as the confidence score (Kadavath et al., 2022).
@@
    Verbalization. This method prompts PoLLMs to express confidence in natural language. We use the prompt from Tian et al. (2023).
@@
    Semantic Entropy (SE). SE (Farquhar et al., 2024; Kuhn et al., 2023) samples multiple responses and computes the entropy over semantic clusters derived from these samples. Following Li et al. (2025), we normalize the SE values to the range $[0,1]$.
@@
    DACA. DACA (Luo et al., 2025), which is designed specifically for the multiple-choice format, uses examples where the PoLLM and base LLM produce the same choice, and then learns a temperature to align the PoLLM’s probabilities with those of the pre-trained model.
Beyond unsupervised baselines, we include Temperature Scaling (TS; Guo et al., 2017) as a supervised reference baseline, which learns a temperature parameter by minimizing the negative log-likelihood with respect to answer correctness. More details are shown in Appendix A.3.
@@ 5.5 Impact of Different PoLLMs (RQ4)
Subsequently, the influence of model size is analyzed using Qwen2.5-7b/14b/32b/72b-instruct. Table 4 shows that BaseCal-Proj consistently yields substantial improvements across all scales. Furthermore, BaseCal-Proj achieves superior performance on larger models, due to the stronger calibration capabilities of larger base LLMs, a trend also observed in previous work (Zhu et al., 2023).
@@ 5.6 Application (RQ5)
Selective classification (Geifman and El-Yaniv, 2017) enables a model to abstain from making predictions when confidence is low, thereby enhancing reliability by trading off coverage for higher accuracy. This mechanism is particularly critical for LLMs in decision-making tasks, where unreliable outputs can lead to severe consequences. In this section, we evaluate the effectiveness of our proposed BaseCal-Proj in the context of selective classification. Figure˜6 compares the accuracy of the vanilla baseline and BaseCal-Proj by varying the confidence threshold from 0.5 to 0.95, where prediction below the threshold is rejected. Experimental results demonstrate that BaseCal-Proj consistently surpasses the vanilla method in accuracy across all confidence thresholds, highlighting its superior capability in identifying reliable predictions.
@@ A.2 Details about LLM-as-a-Judge
To assess the correctness of model outputs in QA tasks, we employ an LLM-as-a-judge to determine whether the generated responses are semantically equivalent to the ground truth, following Tian et al. (2023); Wei et al. (2024). For reproducibility, we utilize the powerful open-source Qwen2.5-14B-Instruct model as our judge. Inspired by the evaluation protocol from Wei et al. (2024), we use the specific prompt detailed in Section˜A.5 to verify response correctness. To validate the quality of the LLM judgment, we conducted a manual audit on 100 randomly sampled instances. Comparative verification against human ground truth reveals a disagreement rate of only 1% with human annotations, confirming the reliability of our automated evaluation process.
@@ A.3 Further Information about Baselines
(1) Vanilla. Several studies have employed the aggregated token probabilities as the vanilla confidence (Orgad et al., 2025; Mahaut et al., 2024; Malinin and Gales, 2021). Following implementation in Orgad et al. (2025), we average the probabilities of all tokens in a response as the confidence.
(2) Verbalization. This method leverages the instruction-following capability of models to express confidence in natural language. Following the implementation of Tian et al. (2023), we employ the following prompt:
@@
(3) P(True). This method follows the prompting strategy introduced by Kadavath et al. (2022), where the LLM is directly queried to assess the correctness of its own output. P(True) also requires an additional forward pass on the PoLLMs after the initial response is generated. Specifically, we construct the following prompt:
@@
(4) Semantic Entropy. As proposed by Kuhn et al. (2023); Farquhar et al. (2024), semantic entropy estimates confidence by sampling multiple responses for the same question and computing their semantic consistency. Following the implementation details recommended by Kuhn et al. (2023), we set the sampling parameters as follows: temperature 0.5, number of samples 10, top_p = 1.0, and top_k = -1. Since this metric was originally designed for hallucination detection, its raw scores do not fall within the $[0,1]$ interval. Therefore, following Li et al. (2025), we normalize the semantic entropy values to the range $[0,1]$ to facilitate the calculation of ECE.
@@
(5) DACA. DACA (Luo et al., 2025) is an unsupervised approach that optimizes a temperature parameter to scale the probability of PoLLM, thereby approximating the probability distribution of the base LLM. This probability-level scaling necessitates that both the base LLM and PoLLM produce the exact same top-1 token for a given input, a constraint theoretically proved and empirically analyzed in Section 3 (pp. 4-5) of Luo et al. (2025). This limitation restricts DACA’s applicability primarily to multiple-choice tasks, as identical token generation is very rare for base LLM and PoLLM in free-form QA. Accordingly, we compare our method with DACA on the MMLU benchmark in the main experiments. We utilize the official implementation (Luo et al., 2025) and optimize DACA using the same training set employed for our proposed method.
(6) Temperature Scaling. We directly adopt the implementation of temperature scaling from Xia et al. (2025). This is a supervised method that optimizes an optimal temperature parameter to rescale probabilities, thereby minimizing the Expected Calibration Error (ECE). The training process requires questions and human-labeled answers to determine response correctness. In our implementation, we utilize the entire training dataset to optimize the temperature.